If you invest $7,000 into a savings account that yields quarterly compound interest with an annual interest rate of 7%, how much will you have in the savings account after 19 years?
Answer #1

Answer: $ 26,165

i = 7 % x 1/4 = 1.75 %

n = 19 x 4 = 76

FVIF = ( 1 + r ) n = ( 1.0175 )76 = 3.7378

Future value of the investment = $ 7,000 x 3.7378 = $ 26,164.6